OpenStudy (anonymous):

can someone help me and show work for this question 2x^4(4x^2 + 3x + 1)

OpenStudy (ivettef365):

just multiply

OpenStudy (ivettef365):

2x^4 (4x^2) = 8x^6

OpenStudy (anonymous):

thats what i am having trouble with because I am not sure how to multiply the coefficient and the exponents

OpenStudy (ivettef365):

then 2x^4 (3x) = 6x^5

OpenStudy (ivettef365):

then 2x^4(1) = 2x^4

OpenStudy (ivettef365):

Therefore you have 8x^6+6x^5+2x^4

OpenStudy (ivettef365):

To multiply exponents you add a^x (a^b) = a^(x+b)

OpenStudy (anonymous):

wow can you help me with a few more please?

OpenStudy (ivettef365):

sure

OpenStudy (anonymous):

how about this one (4x – 3)(2x^2 – 7x + 1)

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

Distributive property

OpenStudy (ivettef365):

That is a long problem, first multiply (4x-3) (2x^2) then by (-7x) then by (1)

OpenStudy (anonymous):

still not sure how to multiply (4x-3) (2x^2)

OpenStudy (ivettef365):

2x4=8 / (x)(x) = x^2 then -3(2) = -6 / and x^2 stays the same, therefore so you have 8x^2-6x^2

OpenStudy (ivettef365):

Do the same process and add all like terms

OpenStudy (ivettef365):

(4x-3)(-7x) you have 4 x 7 - 28 / (x)(x) = x^2 (-3)(-7x) = 21x therefore here you have 28x^2+21x

OpenStudy (ivettef365):

You have then (4x-3) (1) = 4x -3 now add everything

OpenStudy (anonymous):

so it would be 30x^2 + 21x??

OpenStudy (anonymous):

@ivettef365

OpenStudy (anonymous):

@ikarus1991